# The universal multiplicity function: counting halos and voids

### v1

@@ Abstract
We present a novel combination of the excursion-set approach with the peak theory formalism in Lagrangian space and provide accurate predictions for halo and void statistics over a wide range of scales. The set-up is based on an effective moving barrier. Besides deriving the corresponding numerical multiplicity function, we introduce a new analytical formula reaching the percent level agreement with the exact numerical solution obtained via Monte Carlo realizations down to small scales, $\sim 10^{12}h^{-1}{\rm M_{\odot}}$. In the void case, we derive the dependence of the effective moving barrier on the void formation threshold, $\delta_{\rm v}$, by comparison against the Lagrangian void size function measured in the Dark Energy and Massive Neutrinos Universe simulations. We discuss the mapping from Lagrangian to Eulerian space for both halos and voids; adopting the spherical symmetry approximation, we obtain a strong agreement at intermediate and large scales. Finally, using the effective moving barrier, we derive Lagrangian void density profiles accurately matching measurements from cosmological simulations, a major achievement towards using void profiles for precision cosmology with the next generation of galaxy surveys.
@@ 1 Introduction
Our knowledge of the underlying cosmological model of the Universe relies on the analysis of summary statistics of observed data. One standard way to theoretically model these statistics is to compute the distribution of over- and under-densities in the initial matter density field, the so-called Lagrangian space, linearly evolved down to the epoch of interest [^1] Note that in the excursion-set and peak theory frameworks the initial density is intended as the matter density field at a redshift high enough to consider it still in the linear regime.. In this perspective, the two main approaches are the excursion-set formalism (Bond et al., 1991) and the theory of Lagrangian density peaks (Bardeen et al., 1986), which have been widely explored to provide theoretical models for the dark matter (DM) halo mass function (HMF), bias, halo correlation functions (Press & Schechter, 1974; Bardeen et al., 1986; Peacock & Heavens, 1990; Bond et al., 1991; Lacey & Cole, 1993; Mo & White, 1996; Porciani et al., 1998; Catelan et al., 1998; Sheth & Tormen, 1999; Maggiore & Riotto, 2010a; Musso & Sheth, 2012, and references therein), and cosmic void size functions (VSF) (Sheth & van de Weygaert (2004), hereafter SVdW; Jennings et al., 2013; Achitouv et al., 2015).
@@
On the one hand, to model in the initial density field the distribution of regions that would form halos or voids in the evolved Universe, the excursion-set approach considers the density contrast field filtered, i.e. smoothed, on various scales at random positions, which can be described by a random walk (Peacock & Heavens, 1990; Bond et al., 1991). Halos and voids are considered formed when the random walks reach the formation threshold, also called barrier. The typical quantity that can be obtained is the multiplicity function, which is the first-crossing distribution of the forming barrier at a given scale.From this quantity, the HMF, the VSF, and the corresponding local bias expansion can be derived (see Desjacques et al., 2018).
@@
On the other hand, the formalism associated to the peak theory models the progenitors of halos (and voids) using local maxima (minima) in the initial density field, filtered at some scale (Bardeen et al., 1986). However, in this case the link between the number density of peaks and the corresponding multiplicity function is not straightforward. A way to obtain a multiplicity function from peak theory is to combine the statistics of Lagrangian density peaks with the excursion-set (Paranjape & Sheth, 2012).
@@
In this work we explore an effective way to combine the excursion-set framework with peak theory, for both halo and void statistics. In particular, we show that this can be obtained using the standard excursion-set framework with an effective formation barrier, which is a moving barrier that does not correspond to the physical formation threshold, but it can capture the statistical properties of halos and voids. We show that this approach can be used to model the multiplicity function and even higher order statistics, such as the density profile. Our approach is valid for both halos and voids; however, here we mainly focus on voids statistics, as in this case a consistent theoretical modeling is still missing.
@@
This paper is organized as follows. In Sec. 2 we review the building blocks of the excursion-set formalism and excursion-set of peaks, revisiting the halo and void formation within these frameworks. In particular, we propose an effective way to merge the excursion-set with peak theory via an effective moving barrier; in Sec. 3 we discuss the exact numerical solution of our new theoretical model and derive an accurate analytical approximation; in Sec. 4 we discuss how to obtain the effective barrier and derive the one for cosmic voids; in Sec. 5 we introduce the Lagrangian to Eulerian mapping for both halos and voids; in Sec. 6 we apply our model to compute the Lagrangian density profile of voids; finally we draw our conclusions in Sec. 7.
2 Halos and voids in Lagrangian space
@@
The excursion-set approach is based on the concept of modeling the statistical properties of halos or voids in Lagrangian space and mapping them to the fully nonlinearly evolved density field, the Eulerian space.
@@
$\begin{split}\delta(\mathbf{q},R)&=\int\mathrm{d}^{3}x\,W(|\mathbf{x}|,R)\delta(\mathbf{q}+\mathbf{x})\\ &=\int\frac{\mathrm{d}^{3}k}{(2\pi)^{3}}W(kR)\delta(\mathbf{k})e^{-i\mathbf{k\cdot q}}\,.\end{split}$ (1)
Halo and void formation in the excursion-set are described by a thresholding process of the filtered Lagrangian field. A halo with a Lagrangian radius $R$ (and corresponding mass $M(R)$) is formed at the Lagrangian position ${\bf q}$ if $R$ is the maximum smoothing scale at which the filtered density matter field, $\delta({\bf q},R)$, crosses the halo formation threshold (Peacock & Heavens, 1990; Bond et al., 1991). Void formation is analogous to the halo case, with the additional condition that the filtered field that crosses the (negative) formation threshold has not crossed the (positive) halo formation threshold at any larger scale (SVdW). This shows that the treatment of halo and void formation in the excursion-set framework requires the following quantities: formation threshold, filtering function, and position.
@@
A halo is a gravitationally bound, virialised object; for this reason, the threshold for halo formation is chosen as the linear density contrast corresponding to the full collapse in the Eulerian space (Bond et al., 1991). At a first approximation, the dynamics of the halo collapse can be incorporated through a scale-dependent threshold, usually called “moving barrier” (Sheth et al., 2001, hereafter SMT; Sheth & Tormen, 2002; Robertson et al., 2009; Paranjape et al., 2012; Elia et al., 2012; Sheth et al., 2013; Ludlow et al., 2014; Borzyszkowski et al., 2014).
Contrary to the halo case, cosmic voids do not experience any particular event in their evolution, which may characterize their formation: an underdense region starts to grow faster with respect to the background and continues its outward expansion forever (in the single-stream regime). It follows that a map from the Lagrangian to Eulerian density contrast of filtered underdense fluctuations always exists. This condition ensures that the properties (density and size) of evolved voids can always be mapped in the linear theory, and vice versa. In this perspective, the linear void formation threshold $\delta_{\rm v}$ can be any negative value (Pisani et al., 2015; Nadathur & Hotchkiss, 2015; Verza et al., 2019; Contarini et al., 2022), that can be mapped in the corresponding Eulerian one.
@@
The filter function $W(kR)$, entering Eq. (1), is statistically related to the halo and void definitions in Lagrangian space (Bond et al., 1991). In addition, the detailed statistical properties of $\delta({\bf q},R)$ depend on the choice of the specific filter function; the most commonly used ones are the top-hat, Gaussian, and sharp-$k$ filters (see Bardeen et al., 1986, for other physically motivated filter functions). In this work we consider the top-hat filter, which in Fourier space is defined as $W(kR)=3j_{1}(kR)/kR$, where $j_{1}(x)=(\sin x-x\cos x)/x^{2}$ is the spherical Bessel function of order 1. This filter has a clear physical interpretation and a well-defined associated Lagrangian mass, $M_{\rm TH}=4\pi R^{3}\rho_{\rm m}/3$, where $\rho_{\rm m}$ is the mean comoving matter density (see Porciani et al., 1998, for details on other filters).
@@
In the classical formulation of the excursion-set model, the Lagrangian position, ${\bf q}$, of the density contrast field is random, and the corresponding derived statistics for halos and voids are obtained weighting over the entire Lagrangian space. It follows that the corresponding multiplicity function describes the fraction of the Lagrangian volume where the filtered field $\delta({\bf q},R)$ (first) reaches the threshold value. Mapping this quantity to the number density of discrete objects is not straightforward. However, massive halos and voids form and evolve around special positions, such as, respectively, maxima and minima in Lagrangian space (Robertson et al., 2009; Ludlow & Porciani, 2011; Massara & Sheth, 2018). In the literature, it was shown that solving the first crossing distribution problem over all the Lagrangian positions leads to a mismatch between the moving barrier measured in cosmological simulations from the distribution of the Lagrangian patches that will form halos in the evolved universe, and the one obtained by fitting the excursion-set multiplicity function against the measured HMF (SMT; Robertson et al., 2009). To fix this issue, Paranjape & Sheth (2012) proposed an approximated method to derive the multiplicity function from the excursion-set over the subset of maxima rather than over the entire space. They found that at large scales (where the adopted approximations are accurate enough), they can recover the difference between the barriers obtained in the two different ways.
@@
In principle, it is possible to derive the exact solution of various statistics by combining the excursion-set with the peak theory. However, this is complicated both analytically and computationally, and only approximate solutions are practically feasible (Paranjape & Sheth, 2012; Paranjape et al., 2013b). On the other hand, the comparison between the approximated excursion-peak multiplicity function of Paranjape & Sheth (2012) with the (approximated) excursion-set one of Musso & Sheth (2012, hearafter MS) and Sheth & Tormen (1999), implicitly suggests an alternative approach to face this problem: the excursion-peak multiplicity function can be derived from the standard excursion-set one by using an effective moving barrier, which however does not correspond to the physical density contrast of Lagrangian halos and voids (Achitouv et al., 2013). In this work, we show that such an effective barrier contains the statistical information of the multiplicity function and beyond. As a final remark, we underline that in this paper we focus on the void and halo statistics in Lagrangian space; we introduce the mapping from Lagrangian to Eulerian in Sec. 5, nevertheless we leave the full implementation for future work.
@@ 3 The multiplicity function from a generic moving barrier
The stochastic evolution of the Lagrangian density contrast field is formally described by the Langevin equations (Peacock & Heavens, 1990; Bond et al., 1991; Porciani et al., 1998). This field is assumed here to be a Gaussian random field; nevertheless, it is possible to extend the following description to account also for non-Gaussianity (Matarrese et al., 2000; Matarrese & Verde, 2008; Carbone et al., 2008; Maggiore & Riotto, 2010b; Achitouv & Corasaniti, 2012; Musso & Sheth, 2014a). The effect of varying the smoothing scale $R$ can be obtained by differentiating Eq. (1):
$\frac{\partial\delta(R)}{\partial R}=\int\frac{\mathrm{d}^{3}k}{(2\pi)^{3}}\frac{\partial W(kR)}{\partial R}\delta(\mathbf{k})e^{-i\mathbf{k\cdot q}}=\mathcal{Q}(R)\,,$ (2)
where $\delta(R)=\delta(\mathbf{q},R)$. This has the form of the Langevin equation, which shows how an infinitesimal change of the scale $R$ affects the smoothed field as a function of the stochastic force $\mathcal{Q}(R)$. The initial condition of this first-order stochastic differential equation is given by $\delta(R)\rightarrow 0$ as $R\rightarrow\infty$. Since Eq. (2) is linear, the stochastic force is also a Gaussian random field with a vanishing expectation value, $\langle\mathcal{Q}(R)\rangle=0$, and it is uniquely described by its correlation function. Therefore, the Langevin equation is fully described by the equation system (Porciani et al., 1998)
$\begin{cases}\dfrac{\partial\delta(R)}{\partial R}=\displaystyle\int\dfrac{\mathrm{d}^{3}k}{(2\pi)^{3}}\dfrac{\partial W(kR)}{\partial R}\delta(\mathbf{k})e^{-i\mathbf{k\cdot q}}=\mathcal{Q}(R)\\[8.61108pt] \langle\mathcal{Q}(R_{1})\mathcal{Q}(R_{2})\rangle=\displaystyle\int_{0}^{\infty}\dfrac{\mathrm{d}k\,k^{2}}{2\pi^{2}}P(k)\dfrac{\partial W(kR_{1})}{\partial R_{1}}\dfrac{\partial W^{*}(kR_{2})}{\partial R_{2}}\,,\end{cases}$ (3)
@@
The numerical multiplicity function
@@
To numerically obtain the statistical quantities we are interested in, such as the multiplicity function, we solve the first crossing problem over a large number of realizations of the stochastic Eq. (3(Bond et al., 1991; Robertson et al., 2009; Paranjape et al., 2012; Achitouv et al., 2015). The direct solution is numerically expensive, to speed up the computation we exploit the Gaussianity of the density contrast field plus the Cholesky method (Nikakhtar et al., 2018). Let us consider the correlation of the field $\delta(R)$ filtered at different smoothing lengths $R$
$\langle\delta(R_{i})\delta(R_{j})\rangle\equiv C_{ij}=\int\frac{{\rm d}k\,k^{2}}{2\pi^{2}}P(k)W(kR_{i})W^{*}(kR_{j})\,.$ (4)
@@
The matrix ${\bf C}$, with elements $C_{ij}$, is real, symmetric, and positive-definite, so it has a unique decomposition ${\bf C}={\bf LL}^{T}$, where ${\bf L}$ is a lower triangular matrix. It is therefore possible to construct a stochastic numerical realization of $\delta(R)$ as
$\delta(R_{i})=\sum_{j}L_{ij}G_{j}\,,$ (5)
@@
where $G_{j}$ is the $j^{\rm th}$ element of a vector of Gaussian variables, following a distribution with zero mean and unit variance. The multiplicity function is computed as the fraction of realizations, $N_{\times}/N_{\rm tot}$, that first cross the formation barrier, $B(R)$, between $R_{i}$ and $R_{i+1}$
$(R_{i}-R_{i+1})f_{R_{i},R_{i+1}}=N_{\times}[B(R_{i+1})]/N_{\rm tot}.$ (6)
A new analytical multiplicity function

### v2

# The universal multiplicity function: counting haloes and voids

@@ Abstract
We present a novel combination of the excursion-set approach with the peak theory formalism in Lagrangian space and provide accurate predictions for halo and void statistics over a wide range of scales. The set-up is based on an effective moving barrier. Besides deriving the corresponding numerical multiplicity function, we introduce a new analytical formula reaching the percent level agreement with the exact numerical solution obtained via Monte Carlo realisations down to small scales, $\sim 10^{12}h^{-1}{\rm M_{\odot}}$. In the void case, we derive the dependence of the effective moving barrier on the void formation threshold, $\delta_{\rm v}$, by comparison against the Lagrangian void size function measured in the DEMNUni simulations. We discuss the mapping from Lagrangian to Eulerian space for both haloes and voids; adopting the spherical symmetry approximation, we obtain a strong agreement at intermediate and large scales. Finally, using the effective moving barrier, we derive Lagrangian void density profiles accurately matching measurements from cosmological simulations, a major achievement towards using void profiles for precision cosmology with the next generation of galaxy surveys.
@@ 1 Introduction
Our knowledge of the underlying cosmological model of the Universe relies on the analysis of summary statistics of observed data. One standard way to theoretically model these statistics is to compute the distribution of over- and under-densities in the initial matter density field, the so-called Lagrangian space, linearly evolved down to the epoch of interest [^1] Note that in the excursion-set and peak theory frameworks the initial density is intended as the matter density field at a redshift high enough to consider it still in the linear regime.. In this perspective, the two main approaches are the excursion-set formalism [1] and the theory of Lagrangian density peaks [2], which have been widely explored to provide theoretical models for the dark matter (DM) halo mass function (HMF), bias, halo correlation functions [3, 2, 4, 1, 5, 6, 7, 8, 9, 10, 11, and references therein], and cosmic void size functions (VSF) [12], hereafter SVdW[13, 14].
@@
On the one hand, to model in the initial density field the distribution of regions that would form haloes or voids in the evolved Universe, the excursion-set approach considers the density contrast field filtered, i.e. smoothed, on various scales at random positions, which can be described by a random walk [4, 1]. haloes and voids are considered formed when the random walks reach the formation threshold, also called barrier. The typical quantity that can be obtained is the multiplicity function, which is the first-crossing distribution of the forming barrier at a given scale. From this quantity, the HMF, the VSF, and the corresponding local bias expansion can be derived [see 15].
@@
On the other hand, the formalism associated to the peak theory models the progenitors of haloes (and voids) using local maxima (minima) in the initial density field, filtered at some scale [2]. However, in this case the link between the number density of peaks and the corresponding multiplicity function is not straightforward. A way to obtain a multiplicity function from peak theory is to combine the statistics of Lagrangian density peaks with the excursion-set [16].
@@
In this work we explore an effective way to combine the excursion-set framework with peak theory, for both halo and void statistics. In particular, we show that this can be obtained using the standard excursion-set framework with an effective formation barrier, which is a moving barrier that does not correspond to the physical formation threshold, but it can capture the statistical properties of haloes and voids. We show that this approach can be used to model the multiplicity function and even higher order statistics, such as the density profile. Our approach is valid for both haloes and voids; however, here we mainly focus on voids statistics, as in this case a consistent theoretical modelling is still missing.
@@
This paper is organised as follows. In Sec. 2 we review the building blocks of the excursion-set formalism and excursion-set of peaks, revisiting the halo and void formation within these frameworks. In particular, we propose an effective way to merge the excursion-set with peak theory via an effective moving barrier; in Sec. 3 we discuss the exact numerical solution of our new theoretical model and derive an accurate analytical approximation; in Sec. 4 we discuss how to obtain the effective barrier and derive the one for cosmic voids; in Sec. 5 we introduce the Lagrangian to Eulerian mapping for both haloes and voids; in Sec. 6 we apply our model to compute the Lagrangian density profile of voids; finally we draw our conclusions in Sec. 7.
2 haloes and voids in Lagrangian space
@@
The excursion-set approach is based on the concept of modelling the statistical properties of haloes or voids in Lagrangian space and mapping them to the fully nonlinearly evolved density field, the Eulerian space.
@@
$\begin{split}\delta(\mathbf{q},R)&=\int\mathrm{d}^{3}x\,W(|\mathbf{x}|,R)\delta(\mathbf{q}+\mathbf{x})\\ &=\int\frac{\mathrm{d}^{3}k}{(2\pi)^{3}}W(kR)\delta(\mathbf{k})e^{-i\mathbf{k\cdot q}}\,.\end{split}$ (2.1)
Halo and void formation in the excursion-set are described by a thresholding process of the filtered Lagrangian field. A halo with a Lagrangian radius $R$ (and corresponding mass $M(R)$) is formed at the Lagrangian position ${\bf q}$ if $R$ is the maximum smoothing scale at which the filtered density matter field, $\delta({\bf q},R)$, crosses the halo formation threshold [4, 1]. Void formation is analogous to the halo case, with the additional condition that the filtered field that crosses the (negative) formation threshold has not crossed the (positive) halo formation threshold at any larger scale [SVdW]. This shows that the treatment of halo and void formation in the excursion-set framework requires the following quantities: formation threshold, filtering function, and position.
@@
A halo is a gravitationally bound, virialized object; for this reason, the threshold for halo formation is chosen as the linear density contrast corresponding to the full collapse in the Eulerian space [1]. At a first approximation, the dynamics of the halo collapse can be incorporated through a scale-dependent threshold, usually called “moving barrier” [24, hereafter SMT; 17, 18, 19, 20, 21, 22, 23].
Contrary to the halo case, cosmic voids do not experience any particular event in their evolution, which may characterize their formation: an underdense region starts to grow faster with respect to the background and continues its outward expansion forever (in the single-stream regime). It follows that a map from the Lagrangian to Eulerian density contrast of filtered underdense fluctuations always exists. This condition ensures that the properties (density and size) of evolved voids can always be mapped in the linear theory, and vice versa. In this perspective, the linear void formation threshold $\delta_{\rm v}$ can be any negative value [25, 26, 27, 28], that can be mapped in the corresponding Eulerian one.
@@
The filter function $W(kR)$, entering Eq. (2.1), is statistically related to the halo and void definitions in Lagrangian space [1]. In addition, the detailed statistical properties of $\delta({\bf q},R)$ depend on the choice of the specific filter function; the most commonly used ones are the top-hat, Gaussian, and sharp-$k$ filters [see 2, for other physically motivated filter functions]. In this work we consider the top-hat filter, which in Fourier space is defined as $W(kR)=3j_{1}(kR)/kR$, where $j_{1}(x)=(\sin x-x\cos x)/x^{2}$ is the spherical Bessel function of order 1. This filter has a clear physical interpretation and a well-defined associated Lagrangian mass, $M_{\rm TH}=4\pi R^{3}\rho_{\rm m}/3$, where $\rho_{\rm m}$ is the mean comoving matter density [see 7, for details on other filters].
@@
In the classical formulation of the excursion-set model, the Lagrangian position, ${\bf q}$, of the density contrast field is random, and the corresponding derived statistics for haloes and voids are obtained weighting over the entire Lagrangian space. It follows that the corresponding multiplicity function describes the fraction of the Lagrangian volume where the filtered field $\delta({\bf q},R)$ (first) reaches the threshold value. Mapping this quantity to the number density of discrete objects is not straightforward. However, massive haloes and voids form and evolve around special positions, such as, respectively, maxima and minima in Lagrangian space [18, 29, 30]. In the literature, it was shown that solving the first crossing distribution problem over all the Lagrangian positions leads to a mismatch between the moving barrier measured in cosmological simulations from the distribution of the Lagrangian patches that will form haloes in the evolved universe, and the one obtained by fitting the excursion-set multiplicity function against the measured HMF [SMT, 18]. To fix this issue, Paranjape and Sheth [16] proposed an approximated method to derive the multiplicity function from the excursion-set over the subset of maxima rather than over the entire space. They found that at large scales (where the adopted approximations are accurate enough), they can recover the difference between the barriers obtained in the two different ways.
@@
In principle, it is possible to derive the exact solution of various statistics by combining the excursion-set with the peak theory. However, this is complicated both analytically and computationally, and only approximate solutions are practically feasible [16, 31]. On the other hand, the comparison between the approximated excursion-peak multiplicity function of Paranjape and Sheth [16] with the (approximated) excursion-set one of Musso and Sheth [11] (hereafter MS) and SMT, implicitly suggests an alternative approach to face this problem: the excursion-peak multiplicity function can be derived from the standard excursion-set one by using an effective moving barrier, which however does not correspond to the physical density contrast of Lagrangian haloes and voids [32]. In this work, we show that such an effective barrier contains the statistical information of the multiplicity function and beyond. As a final remark, we underline that in this paper we focus on the void and halo statistics in Lagrangian space; we introduce the mapping from Lagrangian to Eulerian in Sec. 5, nevertheless we leave the full implementation for future work.
@@ 3 The multiplicity function from a generic moving barrier
The stochastic evolution of the Lagrangian density contrast field is formally described by the Langevin equations [4, 1, 7]. This field is assumed here to be a Gaussian random field; nevertheless, it is possible to extend the following description to account also for non-Gaussianity [33, 34, 35, 36, 37, 38]. The effect of varying the smoothing scale $R$ can be obtained by differentiating Eq. (2.1):
$\frac{\partial\delta(R)}{\partial R}=\int\frac{\mathrm{d}^{3}k}{(2\pi)^{3}}\frac{\partial W(kR)}{\partial R}\delta(\mathbf{k})e^{-i\mathbf{k\cdot q}}=\mathcal{Q}(R)\,,$ (3.1)
where $\delta(R)=\delta(\mathbf{q},R)$. This has the form of the Langevin equation, which shows how an infinitesimal change of the scale $R$ affects the smoothed field as a function of the stochastic force $\mathcal{Q}(R)$. The initial condition of this first-order stochastic differential equation is given by $\delta(R)\rightarrow 0$ as $R\rightarrow\infty$. Since Eq. (3.1) is linear, the stochastic force is also a Gaussian random field with a vanishing expectation value, $\langle\mathcal{Q}(R)\rangle=0$, and it is uniquely described by its correlation function. Therefore, the Langevin equation is fully described by the equation system [7]
$\begin{cases}\dfrac{\partial\delta(R)}{\partial R}=\displaystyle\int\dfrac{\mathrm{d}^{3}k}{(2\pi)^{3}}\dfrac{\partial W(kR)}{\partial R}\delta(\mathbf{k})e^{-i\mathbf{k\cdot q}}=\mathcal{Q}(R)\\[8.61108pt] \langle\mathcal{Q}(R_{1})\mathcal{Q}(R_{2})\rangle=\displaystyle\int_{0}^{\infty}\dfrac{\mathrm{d}k\,k^{2}}{2\pi^{2}}P(k)\dfrac{\partial W(kR_{1})}{\partial R_{1}}\dfrac{\partial W^{*}(kR_{2})}{\partial R_{2}}\,,\end{cases}$ (3.2)
@@
3.1 The numerical multiplicity function
@@
To numerically obtain the statistical quantities we are interested in, such as the multiplicity function, we solve the first crossing problem over a large number of realisations of the stochastic Eq. (3.2[1, 18, 19, 14]. The direct solution is numerically expensive, to speed up the computation we exploit the Gaussianity of the density contrast field plus the Cholesky method [39]. Let us consider the correlation of the field $\delta(R)$ filtered at different smoothing lengths $R$
$\langle\delta(R_{i})\delta(R_{j})\rangle\equiv C_{ij}=\int\frac{{\rm d}k\,k^{2}}{2\pi^{2}}P(k)W(kR_{i})W^{*}(kR_{j})\,.$ (3.3)
@@
The matrix ${\bf C}$, with elements $C_{ij}$, is real, symmetric, and positive-definite, so it has a unique decomposition ${\bf C}={\bf LL}^{T}$, where ${\bf L}$ is a lower triangular matrix. It is therefore possible to construct a stochastic numerical realisation of $\delta(R)$ as
$\delta(R_{i})=\sum_{j}L_{ij}G_{j}\,,$ (3.4)
@@
where $G_{j}$ is the $j^{\rm th}$ element of a vector of Gaussian variables, following a distribution with zero mean and unit variance. The multiplicity function is computed as the fraction of realisations, $N_{\times}/N_{\rm tot}$, that first cross the formation barrier, $B(R)$, between $R_{i}$ and $R_{i+1}$
$(R_{i}-R_{i+1})f_{R_{i},R_{i+1}}=N_{\times}[B(R_{i+1})]/N_{\rm tot}.$ (3.5)
3.2 A new analytical multiplicity function